# FinDKG: Dynamic Knowledge Graphs with Large Language Models for Detecting Global Trends in Financial Markets

### v1

@@ 4.3. Dynamic knowledge graph learning
(21) $\mathcal{L}=$ $-\sum_{t=1}^{T}\sum_{(s,r,o,t)\in\tilde{\mathcal{G}_{t}}}\bigg{\{}\ \lambda_{1}\log p(t\mid s,r,o,\mathcal{G}_{t})\ +$
(22) $+\lambda_{2}\big{[}\log p(o\mid\mathcal{G}_{t})+\log p(r\mid o,\mathcal{G}_{t})+\log(s\mid r,o,\mathcal{G}_{t})\ +$
(23) $+\log p(s\mid\mathcal{G}_{t})+\log p(r\mid s,\mathcal{G}_{t})+\log(o\mid r,s,\mathcal{G}_{t-1})\big{]}\bigg{\}},$
@@ 5.1. Link prediction on real-world DKGs
We evaluate the performance of the proposed KGTransformer architecture on publicly accessible real-world DKGs used as benchmarks in the literature (Park et al., 2022), alongside the FinDKG introduced as part of this work, described in Section 3.1. Summary statistics about these datasets are described in Table 3. It must be remarked that the only dataset containing meta-entities is FinDKG: therefore, we expect the benefits of KGTransformer to be particularly evident for this dataset. For the other benchmarks, the identity mapping function $\tau(s)=s$ is used, implying that $\mathcal{E}=\mathcal{C}_{\mathcal{E}}$.
@@ 5.3. FinDKG-based thematic investing
Thematic investing is an investment strategy that targets specific themes or trends that are anticipated to influence the future landscape of industries and economies. We demonstrate the utility of FinDKG in estimating corporate exposure to AI, increasingly popular since the launch of OpenAI’s ChatGPT. The objective is to quantitatively measure how closely aligned stock entities are to the prevalence of the AI theme and to generate forward-looking exposure scores.
In an online learning setting, we fit a KGTransformer model within the three-year rolling window FinDKGs at the end of every quarter. At each time $t$, the fitted KGTransformer is used to predict which stock entities are likely to be impacted by AI in the upcoming period $t+1$, corresponding to the quadruple (AI, Impact, ?, $t+1$). Only stocks with a predicted impact likelihood exceeding the average across all entities are retained. This selection forms the basis of a monthly-rebalanced, AI-focused long-only portfolio within the US S&P 500. The portfolio is constructed by using the normalised predicted likelihood scores as the holding weight.
The out-of-sample backtesting results in Table 5 show the efficacy of the FinDKG-based AI portfolio: FinDKG-AI achieves the highest annualized return and Sharpe ratio across all portfolios. The existing AI ETFs lag behind the market benchmark with less return and comparably larger risk. In contrast, the FinDKG AI portfolio outperforms competitors across the evaluation period, with a jump coinciding approximately with the release of OpenAI’s ChatGPT in November 2022, as shown in Figure 6.
@@
FinDKG-AI FinDKG Model-predicted AI Portfolio 39.6% 21.9% 1.810 -18.2%
@@ 6. Conclusion
In this work, we provided three contributions around the use of dynamic knowledge graphs (DKGs) and large language models (LLMs) within financial applications. First, we investigated the performance of fine-tuned open-source LLMs in generating knowledge graphs, proposing the novel open-source Integrated Contextual Knowledge Graph Generator (ICKG) LLM. Next, the ICKG LLM is used to create an open-source dataset from a corpus of financial news articles, called FinDKG. Additionally, we proposed an attention-based architecture called KGTransformer, which incorporates information from meta-entities within the learning process, combining architectures such as HGT (Hu et al., 2020) and EvoKG (Park et al., 2022). Our findings show that the proposed KGTransformer architecture improves the state-of-the-art link prediction performance on two benchmark datasets, and it achieves the best performance with over 10% uplift on FinDKG. Code associated with this work can be found in the GitHub repository xiaohui-victor-li/FinDKG, and an online portal to visualise FinDKG is available at https://xiaohui-victor-li.github.io/FinDKG/.

### v2

@@ 4.3. Dynamic knowledge graph learning
(21) $\mathcal{L}=$ $-\sum_{t=1}^{T}\sum_{(s,r,o,t)\in\tilde{\mathcal{G}_{t}}}\bigg{\{}\ \lambda_{1}\log p(t\mid s,r,o,\mathcal{G}_{t-1})$
$+\lambda_{2}\big{[}\log p(o\mid\mathcal{G}_{t-1})+\log p(r\mid o,\mathcal{G}_{t-1})+\log(s\mid r,o,\mathcal{G}_{t-1})$
(22) $+\log p(s\mid\mathcal{G}_{t-1})+\log p(r\mid s,\mathcal{G}_{t-1})+\log(o\mid r,s,\mathcal{G}_{t-1})\big{]}\bigg{\}},$
@@ 5.1. Link prediction on real-world DKGs
We evaluate the performance of the proposed KGTransformer architecture on publicly accessible real-world DKGs used as benchmarks in the literature (Park et al., 2022), alongside the FinDKG introduced as part of this work, described in Section 3.1. Summary statistics about these datasets are described in Table 3. ICEWS consists in a collection of cooperative or hostile actions between socio-political actors corresponding to individuals, groups, sectors and nation states (Boschee et al., 2015). YAGO dataset is a large, automatically generated knowledge base that combines relational data from Wikipedia and WordNet (Suchanek et al., 2007), whereas WIKI is based on triplets extracted from the Wikidata database (Vrandečić and Krötzsch, 2014; Leblay and Chekol, 2018).
It must be remarked that the only dataset containing meta-entities is FinDKG: therefore, we expect the benefits of KGTransformer to be particularly evident for this dataset. For the other benchmarks, the identity mapping function $\tau(s)=s$ is used, implying that $\mathcal{E}=\mathcal{C}_{\mathcal{E}}$.
@@ 5.3. FinDKG-based thematic investing
Thematic investing is an investment strategy that targets specific themes or trends that are anticipated to influence the future landscape of industries and economies. We demonstrate the utility of FinDKG and link prediction with KGTransformer in estimating corporate exposure to AI, increasingly popular since the launch of OpenAI’s ChatGPT. The objective is to quantitatively measure how closely aligned stock entities are to the prevalence of the AI theme and to generate forward-looking exposure scores.
In an online learning setting, we fit a KGTransformer model within the three-year rolling window FinDKGs at the end of every quarter. At each time $t$ (corresponding to the end of each month), the fitted KGTransformer is used to predict which stock entities are likely to be impacted by AI in the upcoming period $t+1$, corresponding to the quadruple (AI, Impact, ?, $t+1$). Only stocks with a predicted impact likelihood exceeding the average across all entities are retained. This selection forms the basis of a monthly-rebalanced, AI-focused long-only portfolio within the US S&P 500. The portfolio is constructed by using the normalised predicted likelihood scores as the holding weight, which sum up to 100%. We denote this KGTransformer-powered portfolio as FinDKG-AI. We also fit an EvoKG model as an alternative KG learning methodology, using the same FinDKG data and settings, thereby constructing an EvoKG-based AI portfolio as a baseline strategy.
The out-of-sample backtesting results in Table 5 show the efficacy of the FinDKG-based AI portfolio: FinDKG-AI achieves the highest annualized return and Sharpe ratio across all portfolios. The existing AI ETFs lag behind the market benchmark with less return and comparably larger risk. In contrast, the KGTransformer-based FinDKG AI portfolio outperforms competitors across the evaluation period, with a jump coinciding approximately with the release of OpenAI’s ChatGPT in November 2022, as shown in Figure 6. The FinDKG-AI portfolio also outperforms the EvoKG-based strategy, aligning with the better link prediction capabilities of the KGTransformer model architecture shown in previous section.
@@
EvoKG FinDKG EvoKG-predicted Portfolio 23.5% 20.2 % 1.163 -19.1%
FinDKG-AI FinDKG KGTransformer-predicted AI Portfolio 39.6% 21.9% 1.810 -18.2%
@@ 6. Conclusion
In this work, we provided three contributions around the use of dynamic knowledge graphs (DKGs) and large language models (LLMs) within financial applications. First, we investigated the performance of fine-tuned open-source LLMs in generating knowledge graphs, proposing the novel open-source Integrated Contextual Knowledge Graph Generator (ICKG) LLM. Next, the ICKG LLM is used to create an open-source dataset from a corpus of financial news articles, called FinDKG. Additionally, we proposed an attention-based architecture called KGTransformer, which incorporates information from meta-entities within the learning process, combining architectures such as HGT (Hu et al., 2020) and EvoKG (Park et al., 2022).
Our findings show that the proposed KGTransformer architecture improves the state-of-the-art link prediction performance on two benchmark datasets, and it achieves the best performance with over 10% uplift on FinDKG. The generalizability of the ICKG LLM extends beyond the financial news and financial domain, as evidenced by applications in the recent literature adopting similar frameworks (Sarmah et al., 2024; Ouyang et al., 2024). Code associated with this work can be found in the GitHub repository xiaohui-victor-li/FinDKG, and an online portal to visualise FinDKG is available at https://xiaohui-victor-li.github.io/FinDKG/.